OpenStudy (anonymous):

Determine whether the given ordered pair is a solution of the system of equations. Use alphabetically ordered. (5, -10): 3x=25-y 6x=60+3y

OpenStudy (radar):

Substitute 5 for x and -10 for y and see which one is true. Hint. Watch the signs.

OpenStudy (anonymous):

radar is this correct? 3(5)=25-(-10) 15=25+10 15=35 ? 6(5) 60+3(-10) 30=30?

OpenStudy (radar):

Good job jupiter